OpenStudy (anonymous):

if sqrt x is a positive integer such that x^(3/4) =c, then sqrt x=?

OpenStudy (anonymous):

c^(1/3) c^(1/4) c^(1/2) c^(2/3) c

ganeshie8 (ganeshie8):

\(\large x^{(\frac{3}{4})} =c\)

ganeshie8 (ganeshie8):

square both sides and see wat u get

ganeshie8 (ganeshie8):

\( \large (x^{\frac{3}{4}})^2 = c^2 \)

ganeshie8 (ganeshie8):

\( \large x^{\frac{3}{2}} = c^2 \)

ganeshie8 (ganeshie8):

take cube root

ganeshie8 (ganeshie8):

\( \large (x^{\frac{3}{2}})^\frac{1}{3} = (c^2)^\frac{1}{3} \)

ganeshie8 (ganeshie8):

\( \large x^{\frac{1}{2}} = c^\frac{2}{3} \)

ganeshie8 (ganeshie8):

\( \large \sqrt{x} = c^\frac{2}{3} \)

OpenStudy (anonymous):

@ganeshie8 thanks!

ganeshie8 (ganeshie8):

np :)